### v1

# Fast Convergence of $\Phi$-Divergence Along the Unadjusted Langevin Algorithm and Proximal Sampler

Abstract
We study the mixing time of two popular discrete time Markov chains in continuous space, the unadjusted Langevin algorithm and the proximal sampler, which are discretizations of the Langevin dynamics. We extend mixing time analyses for these Markov chains to hold in $\Phi$-divergence. We show that any $\Phi$-divergence arising from a twice-differentiable strictly convex function $\Phi$ converges to 00 exponentially fast along these Markov chains, under the assumption that their stationary distributions satisfies the corresponding $\Phi$-Sobolev inequality. Our rates of convergence are tight and include as special cases popular mixing time regimes, namely the mixing in chi-squared divergence under a Poincaré inequality, and the mixing in relative entropy under a log-Sobolev inequality. Our results follow by bounding the contraction coefficients arising in the appropriate strong data processing inequalities.
1 Introduction
Sampling from a probability distribution is a fundamental task that appears in many fields, including machine learning, statistics, and Bayesian inference [GCSR95, Mac03, RCC99, VT11, JP03]. Suppose we wish to obtain samples from a continuous probability distribution $\nu\propto e^{-f}$ on $\mathbb{R}^{d}$; a common approach is to construct a Markov chain which admits $\nu$ as its invariant or stationary distribution, and then draw samples after some initial burn-in time. A rigorous study of burn-in time [GRS95, Gey11] leads one to analyze the mixing time of a Markov chain, which tracks how quickly the Markov chain converges to its stationary distribution [LPW17]. Different notions of convergence between probability distributions lead to different guarantees, which can be bounded in terms of each other, to yield interesting results and bounds.
A general, useful, and well-studied family of divergences are that of $\Phi$-divergences or $f$-divergences [Csi67], which include many popular divergences such as KL divergence, chi-squared divergence, Total Variation (TV) distance, and squared-Hellinger distance. For any convex function $\Phi:\mathbb{R}_{\geq 0}\to\mathbb{R}$ with $\Phi(1)=0$, the $\Phi$-divergence between probability distributions $\mu$ and $\nu$ such that $\mu\ll\nu$ is defined by
$\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)\coloneqq\mathop{\mathbb{E}}_{\nu}\left[\Phi\left(\frac{\,\mathrm{d}\mu}{\,\mathrm{d}\nu}\right)\right],$ (1)
[…]
1.1 Langevin dynamics
The Langevin dynamics to sample from $\nu\propto e^{-f}$ on $\mathbb{R}^{d}$ is the following SDE:
$\,\mathrm{d}X_{t}=-\nabla f(X_{t})\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\,,$ (2)
where $W_{t}$ is standard Brownian motion on $\mathbb{R}^{d}$. These dynamics admit $\nu$ as the stationary or invariant distribution, and hence are a suitable process to study [BGL14]. The Langevin dynamics being a continuous-time Markov process need to be discretized in time in order to implement. We will focus on two discretizations of the Langevin dynamics – the Unadjusted Langevin Algorithm (ULA) [RT96, RR98, Dal17, CB18, VW23, CEL+22], and the proximal sampler [LST21, CCSW22, KVZ24, KZ24].
The convergence of $\Phi$-divergence along Langevin dynamics is easy to establish (Corollary 1) and the properties of $\Phi$-divergence along continuous-time dynamics are well-studied [Cha04, DL18, AAS15]. We study the convergence of $\Phi$-divergence for discrete-time samplers in Theorems 1 and 2.
1.2 Unadjusted Langevin algorithm
The Unadjusted Langevin Algorithm (ULA) is a simple forward discretization of the Langevin dynamics (2) and is given by the following update
$X_{k+1}=X_{k}-\eta\nabla f(X_{k})+\sqrt{2\eta}Z_{k}\,,$ (3)
where $Z_{k}\sim\mathcal{N}(0,I)$ and $\eta>0$ is the stepsize. As stated in (3), the ULA is biased, meaning that it admits as its stationary distribution $\nu^{\eta}\neq\nu$, which converges to the desired target distribution $\nu$ as $\eta\to 0$ and the ULA (3) recovers the Langevin dynamics (2[RT96]. The ULA has been widely studied, and we now have guarantees on its mixing time in numerous settings – [Dal17] and [CB18] study the mixing time to $\nu$ in KL divergence under $\nu$ being strongly log-concave; [VW23] study mixing to $\nu$ in KL divergence and Rényi divergence under a log-Sobolev inequality and Poincaré inequality respectively, and also to $\nu^{\eta}$ in Rényi divergence under a log-Sobolev inequality assumption on $\nu^{\eta}$[AT23] study mixing to $\nu^{\eta}$ under a strong log-concavity assumption on $\nu$ and in KL divergence, TV distance, Rényi divergence, chi-squared divergence, and squared Hellinger distance; and [CEL+22] study mixing to $\nu$ under modified log-Sobolev inequality assumptions on $\nu$. We refer the reader to [Che24] for a comprehensive overview.
All of the aforementioned works assume smoothness assumptions on $\nu$ (i.e. uniform upper and lower bounds on the Hessian of $f$). While this is common for discrete time analysis, it is a strong assumption, and attempts to go beyond smoothness have had limited success [CDJB20]. In Theorem 1, we show mixing guarantees for ULA in $\Phi$-divergence for smooth target distributions $\nu$ and under the assumption that the stationary distribution $\nu^{\eta}$ of ULA satisfies the corresponding $\Phi$-Sobolev inequality (Definition 1). To the best of our knowledge, we are the first to study mixing time guarantees in any $\Phi$-divergence for ULA. We prove Theorem 1 in Section 3.2.
Theorem 1.
Suppose the stationary distribution of ULA $\nu^{\eta}$ satisfies a $\Phi$-Sobolev inequality with optimal constant $\alpha>0$, and $\nu$ is $L$-smooth for some $0<\alpha\leq L<\infty$. Let $X_{k}\sim\rho_{k}$ evolve following ULA (3) with step size $\eta\leq 1/L$ from $X_{0}\sim\rho_{0}$. Then for all $k\geq 1$,
$\mathsf{D}_{\Phi}(\rho_{k}\mathbin{\|}\nu^{\eta})\leq\left(1+\frac{2\alpha\eta}{(1+\eta L)^{2}}\right)^{-k}\mathsf{D}_{\Phi}(\rho_{0}\mathbin{\|}\nu^{\eta}).$ (4)
As $\eta\to 0$ and $\eta k\to t$, Theorem 1 recovers the correct $\exp(-2\alpha t)$ convergence rate for Langevin dynamics under a $\Phi$-Sobolev inequality (Corollary 1). We show that the rate of Theorem 1 is tight by doing explicit calculations for the case when $\nu$ is a Gaussian in Proposition 1 in Appendix G.1.
A common setting in which $\nu^{\eta}$ satisfies a $\Phi$-Sobolev inequality is when $\nu$ is strongly log-concave (Lemma 6). It is interesting to ask for properties of $\nu^{\eta}$ under weaker assumptions on $\nu$ (see Section 5). For example, similar to how log-Sobolev and Poincaré inequalities are stable to bounded perturbations [HS87], $\Phi$-Sobolev inequalities also enjoy similar properties [Cha04, Section 3.3]. Therefore one can ask if the stationary distribution $\nu^{\eta}$ continues to satisfy a $\Phi$-Sobolev inequality when $\nu$ undergoes a suitable perturbation.
1.3 Proximal sampler
While the simplicity of ULA is appealing, its biased limiting distribution and smoothness assumptions hold it back. The proximal sampler [LST21, CCSW22] addresses both of these shortcomings. Given the close connections between optimization and sampling, along with the modern interpretation of sampling as optimization in the space of measures [JKO98, Wib18], the proximal sampler as introduced in [LST21] can be seen to be the analogue of the proximal point algorithm in optimization [Mar70, Roc76]. The proximal sampler was analyzed in great generality in [CCSW22], who show mixing guarantees in KL divergence and Rényi divergence under various isoperimetric assumptions on $\nu$. In Theorem 2, we show mixing guarantees for the proximal sampler in $\Phi$-divergence under the assumption that $\nu$ satisfies the corresponding $\Phi$-Sobolev inequality (Definition 1). Our results extend the mixing time of the proximal sampler (or any other Langevin based algorithm) to $\Phi$-divergence.
The proximal sampler considers an augmented $(X,Y)$ space $\mathbb{R}^{d}\times\mathbb{R}^{d}$ and performs Gibbs sampling on the joint space. We will use the appropriate superscripts to denote probability distributions on their respective spaces. For example, $\rho^{XY}$ denotes a distribution on the joint space, $\rho^{X}$ refers to a distribution on the $X$ space, and $\rho^{Y\mid X=x}$ denotes a conditional distribution supported on the $Y$ space.
Let the target distribution we wish to sample from be $\nu^{X}\propto e^{-f}$ on $\mathbb{R}^{d}$. The joint target distribution is then defined as:
$\nu^{XY}(x,y)\propto\exp\left(-f(x)-\frac{\|x-y\|^{2}}{2\eta}\right),$ (5)
for step size $\eta>0$. The proximal sampler, initialized from $X_{0}\sim\rho_{0}^{X}$, is the following two-step algorithm.
$\begin{split}&\textbf{Step 1 (forward step)\,:}\text{\,\, Sample \,\,}Y_{k}\mid X_{k}\sim\nu^{Y\mid X=X_{k}}=\mathcal{N}(X_{k},\eta I)\hskip 99.58464pt\\ &\textbf{Step 2 (backward step)\,:}\text{\,\, Sample \,\,}X_{k+1}\mid Y_{k}\sim\nu^{X\mid Y=Y_{k}}\end{split}$ (6)
Note that $\nu^{XY}$ has the desired target distribution $\nu^{X}$ as the $X$ marginal, and that the $Y$ marginal $\nu^{Y}$ is a smoothed version of $\nu^{X}$, i.e. $\nu^{Y}=\nu^{X}*\mathcal{N}(0,\eta I)$. The forward step of the algorithm is easy to implement as it corresponds to drawing a Gaussian random variable. For the backward step, implementation is possible given access to a Restricted Gaussian Oracle (RGO). A RGO is an oracle that, given any $y\in\mathbb{R}^{d}$, outputs a sample from $\nu^{X\mid Y=y}$, i.e. from
$\nu^{X\mid Y}(x\mid y)\propto_{x}\exp\left(-f(x)-\frac{\|x-y\|^{2}}{2\eta}\right).$
Similar to [LST21, CCSW22], we consider an ideal implementation of the proximal sampler where we have exact access to a RGO. Specific cases where the RGO can be implemented efficiently are discussed in [LST21, MFWB22].
More general implementations of the RGO can be done in several different ways, and we mention a basic approach via rejection sampling in Appendix B. Improved implementations of the RGO with minimal assumptions and better complexity is an active area of research. [FYC23] mention an implementation via approximate rejection sampling and [AC24] specify an approach based on Metropolis adjusted Langevin algorithm.
When both steps are implemented exactly (i.e. given access to an exact RGO), the proximal sampler corresponds to Gibbs sampling from stationary distribution $\nu^{XY}$, and the algorithm is therefore unbiased. Step 1 is referred to as the forward step as it corresponds to evolving along the (forward) heat flow, and step 2 is referred to as the backward step is it corresponds to backward heat flow; this perspective is further made clear in Sections 4.1 and 4.2. We now state the main theorem describing the mixing time of the proximal sampler in $\Phi$-divergence. We prove Theorem 2 in Section 4.3.
Theorem 2.
Suppose $\nu^{X}$ satisfies a $\Phi$-Sobolev inequality with optimal constant $\alpha$. Let $X_{k}\sim\rho_{k}^{X}$ evolve along the proximal sampler (6) with step size $\eta>0$ from $X_{0}\sim\rho_{0}^{X}$. Then for all $k\geq 1$,
$\mathsf{D}_{\Phi}(\rho_{k}^{X}\mathbin{\|}\nu^{X})\leq\frac{\mathsf{D}_{\Phi}(\rho_{0}^{X}\mathbin{\|}\nu^{X})}{(1+\alpha\eta)^{2k}}\,.$
We show the tightness of Theorem 2 by doing explicit calculations for the case when $\nu$ is Gaussian in Proposition 2 in Appendix G.2. The setting of Theorem 2 includes popular mixing time results as special cases. For example, it directly implies the mixing in KL divergence under a log-Sobolev inequality, and mixing in chi-squared divergence under a Poincaré inequality. And under a strong log-concavity assumption, it implies mixing in all $\Phi$-divergences. $\Phi$-Sobolev inequalities and $\Phi$-divergences for different choices of $\Phi$ are mentioned in Table 1 in Appendix A.
1.4 Related work
Our main results, Theorems 1 and 2, study the mixing time in $\Phi$-divergence of the ULA (3) and proximal sampler (6) respectively. We give an overview of the algorithms along with prior mixing time results in Sections 1.2 and 1.2, and refer the reader to those sections for the corresponding references. We discuss other related works here, but before proceeding would like to mention that a comprehensive overview of Langevin based samplers can be found in [Che24].
Both the ULA (3) and proximal sampler (6) have the Langevin dynamics (2) as the limiting dynamics as $\eta\to 0$, and the convergence of $\Phi$-divergence along the Langevin dynamics is well-known (see Corollary 1, and also see [Cha04, DL18, AAS15] for a general discussion for diffusions). To study the convergence for discrete-time algorithms, our main tool is that of Strong Data Processing Inequalities (SDPIs) (see Section 2.2). SDPIs are Markov chain-dependent strengthenings of the data processing inequality, and are a fundamental concept in information theory. A comprehensive overview of SDPIs can be found in [PW24, Chapter 33]. Just as data processing inequalities hold in many different metrics and forms, so do SDPIs. [Rag16] study SDPIs in $\Phi$-divergence for discrete-space Markov chains, and [dPCPW17, PW17] study SDPIs in lesser generality than $\Phi$-divergences, but for continuous-space chains and networks.
SDPI inspired techniques have been used in prior works studying the mixing time of Langevin based algorithms, but all of them have been implicit, without connecting back to SDPIs or contraction coefficients (Definition 2), and none have studied $\Phi$-divergences in general. [VW23] use them to study the convergence of ULA to $\nu^{\eta}$ in Rényi divergence, [CCSW22] use them for the proximal sampler, [YFL+23] use them for the proximal sampler on graphs, [KVZ24, KZ24] use them for the constrained proximal sampler. We make the SDPI based approach explicit and use it for $\Phi$-divergence.
Bounding the contraction coefficient (Definition 2) is the key step in showing a Markov chain satisfies a SDPI, and the main idea used to bound the contraction coefficient for Langevin based Markov chains is by taking the time derivative of the divergence along two simultaneous stochastic processes (Lemma 4). This idea traces back to de Bruijn’s identity [Sta59], and similar methods have been used well beyond information theory, for example to study diffusions and diffusion models [Cha04, ABVE23, VW23, KVZ24].
Organization
We go over the necessary background material in Section 2, and then discuss the convergence of $\Phi$-divergence along ULA in Section 3, and along the proximal sampler in Section 4. We summarize and discuss open questions in Section 5 to conclude.
2 Preliminaries
A distribution $\nu\propto e^{-f}$ on $\mathbb{R}^{d}$ with a twice-differentiable potential function $f$ is $\alpha$-strongly log-concave for some $\alpha>0$ if $\alpha I\preceq\nabla^{2}f$, and is $L$-smooth for some $L>0$ if $-LI\preceq\nabla^{2}f\preceq LI$. When $\alpha=0$, we call $\nu$ (weakly) log-concave. Throughout, we take $\Phi\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ to be a twice-differentiable strictly convex function with $\Phi(1)=0$. Whenever we refer to any probability distribution, we always take it to be a member of $\mathcal{P}_{2,ac}(\mathbb{R}^{d})$, i.e. the set of probability distributions on $\mathbb{R}^{d}$ which are absolutely continuous with respect to Lebesgue measure and have finite second moment. We also abuse notation and refer to distributions via their densities with respect to Lebesgue measure. On occasion, we also denote $\mathcal{N}_{t}$ as shorthand for $\mathcal{N}(0,tI)$ where $t>0$ and $\mathcal{N}(\mu,\Sigma)$ refers to a Gaussian distribution with mean $\mu\in\mathbb{R}^{d}$ and positive-definite covariance matrix $\Sigma\in\mathbb{R}^{d\times d}$.
2.1 $\Phi$-divergences, $\Phi$-Sobolev inequalities, and their properties
We now define $\Phi$-Sobolev inequalities, which include as special cases popular isoperimetric inequalities such as log-Sobolev inequalities and Poincaré inequalities.
Definition 1.
A probability distribution $\nu$ satisfies a $\Phi$-Sobolev inequality $(\Phi\mathsf{SI})$ with constant $\alpha>0$ if for all probability distributions $\mu\ll\nu$ , we have
$2\alpha\,\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)\leq\mathsf{FI}_{\Phi}(\mu\mathbin{\|}\nu),$ (7)
where $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)$ is defined in (1) and
$\mathsf{FI}_{\Phi}(\mu\mathbin{\|}\nu)$ $\coloneqq\mathop{\mathbb{E}}_{\nu}\left[\left\|\nabla\frac{\,\mathrm{d}\mu}{\,\mathrm{d}\nu}\right\|^{2}\Phi^{\prime\prime}\left(\frac{\,\mathrm{d}\mu}{\,\mathrm{d}\nu}\right)\right].$
We define the $\Phi$-Sobolev constant of $\nu$ to be the optimal (largest) constant $\alpha$ such that the above inequality holds:
$\alpha_{\Phi\mathsf{SI}}(\nu):=\inf_{\mu}\frac{\mathsf{FI}_{\Phi}(\mu\mathbin{\|}\nu)}{2\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)}$ (8)
where the infimum is taken over all probability distributions $\mu$ with $0<\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)<\infty$.
The inequality (7) is equivalent to saying that for all smooth functions $g:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ with $\mathop{\mathbb{E}}_{\nu}[g]=1$,
$2\alpha\,\mathsf{Ent}_{\Phi}^{\nu}(g)\leq\mathcal{E}_{\Phi}^{\nu}(g)\,,$ (9)
where
$\mathsf{Ent}_{\Phi}^{\nu}(g)\coloneqq\mathop{\mathbb{E}}_{\nu}[\Phi(g)]-\Phi(\mathop{\mathbb{E}}_{\nu}[g])\text{ \,\,and\,\, }\mathcal{E}_{\Phi}^{\nu}(g):=\mathop{\mathbb{E}}_{\nu}\left[\|\nabla g\|^{2}\Phi^{\prime\prime}(g)\right].$
To see this, take $g$ to be the density function of $\mu$ with respect to $\nu$.
For $\Phi(x)=x\log x$, the $\Phi$-divergence is the KL divergence and the $\Phi$-Sobolev inequality is the log-Sobolev inequality. For $\Phi(x)=(x-1)^{2}$, the $\Phi$-divergence is the chi-squared divergence and the $\Phi$-Sobolev inequality is the Poincaré inequality. Further examples are mentioned in Table 1 in Appendix A.
Core to our SDPI based proof strategy will be analyzing the change of the $\Phi$-Sobolev constant along various operations such as convolution and pushforward. These operations arise by interpreting the Markov chains (3) and (6) as updates in the space of distributions, and will be described in Sections 3 and 4 respectively. The evolution of the $\Phi$-Sobolev constant along these operations is classical [Cha04].
The following lemma tells us how the $\Phi\mathsf{SI}$ constant evolves along a Lipschitz pushforward map.
Lemma 1.
[Cha04, Remark 7] Assume $\nu$ satisfy $\Phi$-Sobolev inequality with optimal constant $\alpha_{\Phi\mathsf{SI}}(\nu)$. Let $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a $\gamma$-Lipschitz map. Then, the pushforward $\tilde{\nu}=T_{\#}\nu$ satisfies $\Phi$-Sobolev inequality with optimal constant
$\alpha_{\Phi\mathsf{SI}}(\tilde{\nu})\geq\frac{\alpha_{\Phi\mathsf{SI}}(\nu)}{\gamma^{2}}.$
The next lemma describes the change of the $\Phi\mathsf{SI}$ constant after convolution.
Lemma 2.
[Cha04, Corollary 3.1] Assume $\mu$ and $\nu$ satisfy $\Phi\mathsf{SI}$ with optimal constants $\alpha_{\Phi\mathsf{SI}}(\mu)$ and $\alpha_{\Phi\mathsf{SI}}(\nu)$, respectively. Then the convolution $\mu\ast\nu$ satisfies $\Phi\mathsf{SI}$ with constant
$\frac{1}{\alpha_{\Phi\mathsf{SI}}(\mu*\nu)}\leq\frac{1}{\alpha_{\Phi\mathsf{SI}}(\mu)}+\frac{1}{\alpha_{\Phi\mathsf{SI}}(\nu)}.$
The following lemma tells us that when $\nu$ is $\alpha$-strongly log-concave, it also satisfies $\Phi$-Sobolev inequality with the same constant.
Lemma 3.
[Cha04, Corollary 2.1] If $\nu$ is $\alpha$-strongly log-concave for some $\alpha>0$, then $\nu$ satisfies $\Phi\mathsf{SI}$ with constant
$\alpha_{\Phi\mathsf{SI}}(\nu)\geq\alpha.$
We conclude this subsection by describing the rate of change of $\Phi$-divergence along simultaneous evolutions of the same SDE. The following lemma will be crucial in the SDPI based approach. We prove Lemma 4 in Appendix C.
Lemma 4.
Suppose $X_{t}\sim\mu_{t}$ and $X_{t}\sim\nu_{t}$ with initial conditions $\mu_{0}$ and $\nu_{0}$ are two solutions of the following SDE:
$\,\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\sqrt{2c}\,\mathrm{d}W_{t}\,,$ (10)
where $b_{t}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a time-varying drift function, $c$ is a positive constant, and $W_{t}$ is standard Brownian motion on $\mathbb{R}^{d}$. Then,
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})=-c\,\mathsf{FI}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t}).$
As mentioned in Section 1.4, when (10) is taken to be Brownian motion (i.e. the drift $b_{t}=0$), $\nu_{t}$ is fixed to be the Lebesgue measure, and $\Phi(x)=x\log x$ , Lemma 4 corresponds to de Bruijn’s identity [Sta59]. The SDE (10) is more general than the Langevin dynamics (2) as it includes a time-varying drift function; but this is required to study the proximal sampler. As we will see in Section 4, the backward step of the proximal sampler corresponds to an SDE with time-varying drift.
As an easy consequence of Lemma 4, we have the exponential convergence of $\Phi$-divergence for the Langevin dynamics under $\nu$ satisfying a $\Phi$-Sobolev inequality.
Corollary 1.
Suppose $X_{t}\sim\rho_{t}$ evolves along the Langevin dynamics (2) to sample from $\nu\propto e^{-f}$, and let $\nu$ satisfy a $\Phi$-Sobolev inequality with optimal constant $\alpha>0$. Then,
$\mathsf{D}_{\Phi}(\rho_{t}\mathbin{\|}\nu)\leq e^{-2\alpha t}\mathsf{D}_{\Phi}(\rho_{0}\mathbin{\|}\nu).$
Proof.
Applying Lemma 4 (with $\mu_{t}=\rho_{t}$ and $\nu_{t}=\nu$) along with the $\Phi$-Sobolev inequality of $\nu$ yields,
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu)=\mathsf{FI}_{\Phi}(\mu_{t}\mathbin{\|}\nu)\leq-2\alpha\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu).$
Integrating the differential inequality from 00 to $t$ using Grönwall’s lemma completes the proof.
2.2 Strong data processing inequalities
We now discuss Strong Data Processing Inequalities (SDPIs), which will be our primary proof strategy. For any $\Phi$-divergence, the data processing inequality states that for any two distributions $\mu$ and $\nu$, and for any Markov kernel $\bm{\mathrm{Q}}$, $\mathsf{D}_{\Phi}(\mu\bm{\mathrm{Q}}\mathbin{\|}\nu\bm{\mathrm{Q}})\leq\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)$, provided all quantities are well-defined [^1] For an example of the Markov kernel notation, and it’s behaviour as on operator on distributions, see (14) and the subsequent text.. Strong data processing inequalities check if the inequality is strict, and quantify the decrease [PW24, PW17, Rag16]. They do so by fixing the second input distribution (i.e. $\nu$) and then varying the first distribution (i.e. $\mu$), to see the worst case ratio $\frac{\mathsf{D}_{\Phi}(\mu\bm{\mathrm{Q}}\mathbin{\|}\nu\bm{\mathrm{Q}})}{\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)}$. The ratio is called the contraction coefficient, and when it is strictly less than $1$ for all valid distributions $\mu$ , we say that $(\bm{\mathrm{Q}},\nu)$ satisfies a strong data processing inequality.
Next, we define the contraction coefficient in Definition 2 and define SDPIs in $\Phi$-divergence in Definition 3.
Definition 2.
Let $\nu$ be a probability distribution, $\bm{\mathrm{Q}}$ be a Markov kernel, and $\mathsf{D}_{\Phi}$ be a $\Phi$-divergence. Then the contraction coefficient $\varepsilon_{\mathsf{D}_{\Phi}}$ is defined as follows
[…]
Definition 3.
Let $\nu$ be a probability distribution, $\bm{\mathrm{Q}}$ be a Markov kernel, and $\mathsf{D}_{\Phi}$ be a $\Phi$-divergence. Further define $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{Q}},\nu)$ as in (11). Then we say that $(\bm{\mathrm{Q}},\nu)$ satisfies a strong data processing inequality in $\Phi$-divergence when $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{Q}},\nu)<1$ In particular, we have the following strengthening of the data processing inequality in $\Phi$-divergence,
$\mathsf{D}_{\Phi}(\mu\bm{\mathrm{Q}}\mathbin{\|}\nu\bm{\mathrm{Q}})\leq\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{Q}},\nu)\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)\,,$ (12)
where $\mu$ is any distribution such that $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)<\infty$
As a direct consequence of Definition 3, we can see that if $\nu$ is invariant for $\bm{\mathrm{Q}}$, then for any $\mu$,
$\mathsf{D}_{\Phi}(\mu\bm{\mathrm{Q}}^{k}\mathbin{\|}\nu)\leq\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{Q}},\nu)^{k}\,\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu).$ (13)
Hence if we desire to show a mixing time result for Markov chain $\bm{\mathrm{Q}}$ with invariant distribution $\nu$, we require a bound on $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{Q}},\nu)$ which is strictly less than $1$. Relationships between the contraction coefficients in various divergences allow one to bound the coefficient of interest by a coefficient that is convenient to bound [SV16], but we will bound the contraction coefficient in the specific $\Phi$-divergence directly.
3 Convergence Along ULA
Recall the ULA update (3) for sampling from $\nu\propto e^{-f}$. Denoting $\rho_{k}=\mathsf{law}(X_{k})$ , the update (3) can be seen as the following update in law
$\rho_{k+1}=(\mathsf{id}-\eta\nabla f)_{\#}\rho_{k}*\mathcal{N}(0,2\eta I)\,.$ (14)
Denote the Markov kernel of ULA as $\bm{\mathrm{P}}$ (i.e. $\rho_{k+1}=\rho_{k}\bm{\mathrm{P}}=\rho_{0}\bm{\mathrm{P}}^{k+1}$) and $F(x)=x-\eta\nabla f(x)$. Recall from Section 1.2 that ULA has a biased stationary distribution $\nu^{\eta}$. Then as per (13), in order to obtain a mixing time result, our goal is to bound $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}},\nu^{\eta})$. We now provide a proof outline for this.
3.1 Proof outline
We wish to bound $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}},\nu^{\eta})$. To that end, let $\mu$ be an arbitrary probability distribution and recall by Definition 2 that we want to control
$\frac{\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}\mathbin{\|}\nu^{\eta}\bm{\mathrm{P}})}{\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{\eta})}=\frac{\mathsf{D}_{\Phi}(F_{\#}\mu*\mathcal{N}(0,2\eta I)\mathbin{\|}F_{\#}\nu^{\eta}*\mathcal{N}(0,2\eta I))}{\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{\eta})}.$
Suppose that $F$ is a bijective map. Then, as $\Phi$-divergences are invariant with respect to simultaneous bijective maps [^2] This can be seen as a consequence of data processing inequality, i.e. for any $\mu$ and $\rho$ and bijective map $F$, $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\rho)\geq\mathsf{D}_{\Phi}(F_{\#}\mu\mathbin{\|}F_{\#}\rho)\geq\mathsf{D}_{\Phi}(F^{-1}_{\#}F_{\#}\mu\mathbin{\|}F^{-1}_{\#}F_{\#}\rho)=\mathsf{D}_{\Phi}(\mu\mathbin{\|}\rho)$, hence all the inequalities must be equalities., $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{\eta})=\mathsf{D}_{\Phi}(F_{\#}\mu\mathbin{\|}F_{\#}\nu^{\eta})$. Therefore, under the assumption that $F$ is bijective, the quantity we wish to bound is
$\frac{\mathsf{D}_{\Phi}(F_{\#}\mu*\mathcal{N}(0,2\eta I)\mathbin{\|}F_{\#}\nu^{\eta}*\mathcal{N}(0,2\eta I))}{\mathsf{D}_{\Phi}(F_{\#}\mu\mathbin{\|}F_{\#}\nu^{\eta})}.$
Denoting $F_{\#}\mu*\mathcal{N}(0,tI)=\mu_{t}$ and $F_{\#}\nu^{\eta}*\mathcal{N}(0,tI)=\nu_{t}$, we want to bound
$\frac{\mathsf{D}_{\Phi}(\mu_{2\eta}\mathbin{\|}\nu_{2\eta})}{\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})}.$
As convolving with a Gaussian can be viewed as the solution to the heat equation, this quantity can be bounded as a consequence of Lemma 4 (for an appropriate SDE, i.e. for (10) with $b_{t}=0$ and $c=\frac{1}{2}$) under the assumption that $\nu_{t}$ satisfies a $\Phi$-Sobolev inequality for $t\in[0,2\eta]$. The final expression obtained in independent of $\mu$, and therefore this provides a bound on $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}},\nu^{\eta})$.
This is the outline we follow and the assumptions stated in Theorem 1 are to ensure that the assumptions mentioned in the proof sketch go through.
3.2 Proof of Theorem 1
In this section we prove Theorem 1. The key lemma for doing so is the following. The proof of Lemma 5 follows the outline mentioned in Section 3.1 and is in Appendix D.
Lemma 5.
Let $\bm{\mathrm{P}}$ denote the ULA Markov kernel with update (14). Suppose $\nu^{\eta}$ satisfies a $\Phi$-Sobolev inequality with optimal constant $\alpha$, $\nu$ is $L$-smooth, and $\eta\leq\frac{1}{L}$. Then,
$\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}},\nu^{\eta})\leq\frac{(1+\eta L)^{2}}{(1+\eta L)^{2}+2\alpha\eta}\,.$
Lemma 5 gives a bound on the contraction coefficient for ULA, with which we can prove Theorem 1.
Proof of Theorem 1.
By an easy application of Definition 3 and by noting that $\nu^{\eta}$ is stationary for $\bm{\mathrm{P}}$, we have that,
$\mathsf{D}_{\Phi}(\rho_{k}\mathbin{\|}\nu^{\eta})\leq\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}},\nu^{\eta})^{k}\mathsf{D}_{\Phi}(\rho_{0}\mathbin{\|}\nu^{\eta})\,.$
Using Lemma 5, we get that,
$\mathsf{D}_{\Phi}(\rho_{k}\mathbin{\|}\nu^{\eta})$ $\leq\left(\frac{(1+\eta L)^{2}}{(1+\eta L)^{2}+2\alpha\eta}\right)^{k}\mathsf{D}_{\Phi}(\rho_{0}\mathbin{\|}\nu^{\eta})=\left(1+\frac{2\alpha\eta}{(1+\eta L)^{2}}\right)^{-k}\mathsf{D}_{\Phi}(\rho_{0}\mathbin{\|}\nu^{\eta})\,.$
Explicit calculations for the Ornstein-Uhlenbeck process implying the tightness of Theorem 1 can be found in Proposition 1 in Appendix G.1.
3.3 Property of the biased limit
We conclude this section by showing that the biased limit satisfies a $\Phi$-Sobolev inequality under strong log-concavity of $\nu$. We prove Lemma 6 in Appendix E.
Lemma 6.
Suppose $\nu\propto e^{-f}$ is $\alpha$-strongly log-concave and $L$-smooth. Consider the ULA (3) to sample from $\nu$ with $\eta\leq\frac{1}{L}$. Then the biased limit $\nu^{\eta}$ satisfies:
$\alpha_{\Phi\mathsf{SI}}(\nu^{\eta})\geq\frac{\alpha(2-\alpha\eta)}{2}\geq\frac{\alpha}{2}\,.$
4 Convergence Along Proximal Sampler
Recall the proximal sampler from Section 1.3 with update given by (6), with the forward and backward step. We will denote the proximal sampler as $\bm{\mathrm{P}}_{\textnormal{\!prox}}=\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-}$ where $\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}$ corresponds to the forward step and $\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-}$ to the backward step. Each step of the proximal sampler is an update on $\mathbb{R}^{d}$, and so this perspective via composition of Markov kernels is valid. In terms of notation, we have $\mathsf{law}(X_{k})=\rho_{k}^{X}$, $\mathsf{law}(Y_{k})=\rho_{k}^{Y}$, and therefore $\rho^{X}_{k}\,\bm{\mathrm{P}}_{\textnormal{\!prox}}=\rho^{X}_{k+1}$ , $\rho^{X}_{k}\,\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}=\rho^{Y}_{k}$ , and so on. When clear from context, we will also denote $\bm{\mathrm{P}}_{\textnormal{\!prox}}$ as $\bm{\mathrm{P}}$, $\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}$ as $\bm{\mathrm{P}}^{+}$, and $\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-}$ as $\bm{\mathrm{P}}^{-}$.
As mentioned in Section 1.3, we know that $\nu^{X}$ is stationary for the proximal sampler, and from (13), our goal is to then bound the contraction coefficient $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}},\nu^{X})$. The following lemma shows that this can be bounded by the product of the contraction coefficients of the forward and backward steps.
Lemma 7.
Let $\bm{\mathrm{P}}_{\textnormal{\!prox}}=\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-}$ denote the proximal sampler (6) with joint stationary distribution $\nu^{XY}$. Then,
$\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}},\nu^{X})\leq\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+},\nu^{X})\,\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-},\nu^{Y})\,.$
Proof.
By Definition 2,
$\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}},\nu^{X})=\sup_{\mu}\frac{\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}\mathbin{\|}\nu^{X}\bm{\mathrm{P}})}{\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{X})}$ $=\sup_{\mu}\left[\frac{\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}^{+}\mathbin{\|}\nu^{X}\bm{\mathrm{P}}^{+})}{\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{X})}\times\frac{\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}^{+}\bm{\mathrm{P}}^{-}\mathbin{\|}\nu^{X}\bm{\mathrm{P}}^{+}\bm{\mathrm{P}}^{-})}{\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}^{+}\mathbin{\|}\nu^{X}\bm{\mathrm{P}}^{+})}\right]$
$\leq\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+},\nu^{X})\,\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-},\nu^{X}\bm{\mathrm{P}}^{+})\,.$
The claim follows by noting that $\nu^{Y}=\nu^{X}\bm{\mathrm{P}}^{+}$ (which is further made clear in Section 4.1). ∎
In light of Lemma 7, we will bound each of the contraction coefficients separately and get our convergence result. As each contraction coefficient can at most equal $1$ (as a consequence of data processing inequality), this perspective implies that showing that either coefficient is strictly less than $1$ yields a convergence guarantee in $\Phi$-divergence.
4.1 Forward step
Without loss of generality consider $k=0$, and recall the forward step of the proximal sampler (6), where $X_{0}\sim\rho_{0}^{X}$ and $Y_{0}\mid X_{0}\sim\mathcal{N}(X_{0},\eta I)$. In order to understand the forward step, we wish to understand $\rho_{0}^{Y}$. For any $y\in\mathbb{R}^{d}$, $\rho_{0}^{Y}(y)=\int\nu^{Y\mid X}(y\mid x)\rho_{0}^{X}(x)\,\mathrm{d}x$, and therefore $\rho_{0}^{Y}=\rho_{0}^{X}*\mathcal{N}(0,\eta I)$. This implies that $\bm{\mathrm{P}}^{+}$ corresponds to evolving along the heat flow (i.e. $\,\mathrm{d}X_{t}=\,\mathrm{d}W_{t}$) for time $\eta$. And so, to get a control on $\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+},\nu^{X})$, we can use Lemma 4 with $b_{t}=0$ and $c=\frac{1}{2}$, just like was done for ULA (discussed in Section 3). This leads to the following guarantee on the contraction coefficient. We prove Lemma 8 in Appendix F.1.
Lemma 8.
Let $\bm{\mathrm{P}}_{\textnormal{\!prox}}=\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-}$ denote the proximal sampler (6) with step size $\eta>0$, to sample from $\nu^{X}$ where $\nu^{X}$ satisfies a $\Phi$-Sobolev inequality with optimal constant $\alpha>0$. Then,
$\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+},\nu^{X})\leq\frac{1}{1+\alpha\eta}.$
4.2 Backward step
We now focus on the backward step of the proximal sampler. For the forward step, we were able to relate it to an SDE ($\,\mathrm{d}X_{t}=\,\mathrm{d}W_{t}$) by observing how it acts on distributions (i.e. on $\rho_{0}^{X}$, and seeing that $\rho_{0}^{Y}=\rho_{0}^{X}*\mathcal{N}(0,\eta I)$). For the backward step, this approach is not as clear. Indeed, writing $\rho^{X}_{1}(x)=\int\nu^{X\mid Y}(x\mid y)\rho_{0}^{Y}(y)\,\mathrm{d}y$ does not immediately yield an SDE interpretation. But by step 2 of the proximal sampler (6), we are looking for an SDE such that when initialized at $y\in\mathbb{R}^{d}$, has marginal law $\nu^{X\mid Y=y}$ at time $\eta$. And therefore in general, when initialized at $\nu^{Y}$, it will yield $\nu^{X}$, and when started at $\rho_{0}^{Y}$, will output $\rho^{X}_{1}$. It will be helpful to think of $\nu^{X}$ as being obtained by the time reversal of $\nu^{Y}$.
For the forward step, we have that $\,\mathrm{d}X_{t}=\,\mathrm{d}W_{t}$ where if $X_{0}\sim\nu^{X}$, then $X_{\eta}\sim\nu^{Y}$. The time reversal of this SDE is called the backward heat flow and is given by
$\,\mathrm{d}Y_{t}=\nabla\log(\nu^{X}*\mathcal{N}_{\eta-t})(Y_{t})\,\mathrm{d}t+\,\mathrm{d}W_{t}.$ (15)
By construction, if we start (15) from $\nu^{Y}$, then for any $t\in[0,\eta]$, the marginal law along (15) is the same as $\nu^{X}*\mathcal{N}_{\eta-t}$. Such a reverse SDE construction is popular in diffusion models [CCL+23], and details regarding it can be found in [Föl05, CCGL23]. Rigorous connections between (15) and the proximal sampler can be found in [Che24, Chapter 8.3] and [CCSW22, Appendix A.1.2].
The following lemma describes the contraction coefficient for the backward step. It is proven in Appendix F.2.
Lemma 9.
Let $\bm{\mathrm{P}}_{\textnormal{\!prox}}=\bm{\mathrm{P}}_{\textnormal{\!prox}}^{+}\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-}$ denote the proximal sampler (6) with step size $\eta>0$, to sample from $\nu^{X}$ where $\nu^{X}$ satisfies a $\Phi$-Sobolev inequality with optimal constant $\alpha>0$. Then,
$\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}}^{-},\nu^{Y})\leq\frac{1}{1+\alpha\eta}.$
The reason that both Lemmas 8 and 9 have the same result is for two reasons. First, it is because Lemma 4 does not depend on the drift $b_{t}$. Second, it is because the marginals of the forward and backward heat flow are in correspondence (via the fact that if we start (15) from $\nu^{Y}$, then for any $t\in[0,\eta]$, the marginal law along (15) is the same as $\nu^{X}*\mathcal{N}_{\eta-t}$).
4.3 Proof of Theorem 2
Equipped with Lemmas 8 and 9, the proof of Theorem 2 follows easily.
Proof of Theorem 2.
By repeated application of Definition 3 and by the fact that $\nu^{X}$ is stationary for the proximal sampler, we have that,
$\mathsf{D}_{\Phi}(\rho_{k}^{X}\mathbin{\|}\nu^{X})\leq\varepsilon_{\mathsf{D}_{\Phi}}(\bm{\mathrm{P}}_{\textnormal{\!prox}},\nu^{X})^{k}\,\mathsf{D}_{\Phi}(\rho_{0}^{X}\mathbin{\|}\nu^{X}).$
Further using Lemmas 78, and 9, we get,
$\mathsf{D}_{\Phi}(\rho_{k}^{X}\mathbin{\|}\nu^{X})\leq\frac{\mathsf{D}_{\Phi}(\rho_{0}^{X}\mathbin{\|}\nu^{X})}{(1+\alpha\eta)^{2k}}.$
Explicit calculations for the Ornstein-Uhlenbeck process implying the tightness of Theorem 2 can be found in Proposition 2 in Appendix G.2.
5 Discussion
We study the mixing time in $\Phi$-divergence for two popular discretizations of the Langevin dynamics, namely the unadjusted Langevin algorithm and the proximal sampler. Our results show mixing to the stationary distributions of these Markov chains under the stationary distributions satisfying $\Phi$-Sobolev inequalities. As the proximal sampler is unbiased, this implies mixing to the target distribution $\nu$ of interest. However, this is not the case for the ULA, where our mixing guarantees are to the biased limit $\nu^{\eta}$. While $\nu^{\eta}$ can be shown to satisfy a $\Phi$-Sobolev inequality under strong log-concavity of $\nu$ (Lemma 6), it is interesting to study what can hold under weaker assumptions on $\nu$. For example one can ask if the fact that $\nu^{\eta}$ satisfies a $\Phi$-Sobolev inequality is closed under certain perturbations of $\nu$. Alternatively, one can also ask for mixing time results to $\nu$ in $\Phi$-divergence directly, and can also pose all of these questions for other Markov chains such as Hamiltonian Monte Carlo and the underdamped Langevin algorithm.
Our results are based on strong data processing inequalities in $\Phi$-divergence and proceed by bounding the respective contraction coefficients. However, different forms of (strong) data processing inequalities exist, namely in terms of mutual information and $\Phi$-mutual information. It is interesting to ask if our approaches can extend to these other forms of strong data processing inequalities, to yield for example, the convergence of the mutual information functional. Additionally, we require the function $\Phi$ to be twice differentiable, which prohibits many popular $\Phi$-divergences such as total variation distance and “hockey-stick” divergences [SV16]. These divergences are popular in differential privacy applications [ADC20] and it would therefore be instructive to study extensions to such divergences directly.
Finally, our results imply mixing in all (twice-differentiable) $\Phi$-divergences under strongly log-concave target distributions. It is interesting to ask if mixing bounds can be obtained in a variety of $\Phi$-divergences under weaker absolute isoperimetric assumptions on $\nu$ (such as a log-Sobolev inequality, or in other words, a $\Phi$-Sobolev inequality with $\Phi(x)=x\log x$).
Acknowledgements
We thank Varun Jog and Sinho Chewi for helpful comments and references.
Appendix A Examples of $\Phi$-divergences
$\Phi(x)$ $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)$ $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu)$ name $\Phi$-Sobolev inequality
[…]
$(x-1)^{2}$bigintsss$\bigintsss\frac{(\,\mathrm{d}\mu-\,\mathrm{d}\nu)^{2}}{\,\mathrm{d}\nu}$ chi-squared divergence Poincaré inequality
$(\sqrt{x}-1)^{2}$bigintsss$\bigintsss(\sqrt{\,\mathrm{d}\mu}-\sqrt{\,\mathrm{d}\nu})^{2}$ squared Hellinger distance
[…]
Table 1: Common $\Phi$ functions along with corresponding $\Phi$-divergences (1) and $\Phi$-Sobolev inequalities (Definition 1).
Appendix B Restricted Gaussian Oracle
Here we show a basic implementation of the RGO via rejection sampling. As mentioned in Section 1.3, enhanced implementations of the RGO under weaker assumptions have been an active area of research and we refer the reader to [FYC23] and the references therein.
Recall from Section 1.3 the conditional distribution the RGO seeks to sample from:
$\nu^{X\mid Y}(x\mid y)\propto_{x}\exp\left(-f(x)-\frac{\|x-y\|^{2}}{2\eta}\right).$
Define $g_{y}(x)\coloneqq f(x)+\frac{\|x-y\|^{2}}{2\eta}$ so that for any fixed $y\in\mathbb{R}^{d}$, the target distribution for the RGO is $\tilde{\nu}_{y}(x)\propto\exp(-g_{y}(x))$. Suppose the potential function $f$ is $L$-smooth and that $\eta<\frac{1}{L}$. In this case, $\tilde{\nu}_{y}$ is strongly log-concave with condition number $\frac{1+L\eta}{1-L\eta}$ and the RGO can be implemented efficiently via rejection sampling.
Suppose $\pi\propto e^{-V}$ is $\beta$-strongly log-concave and $M$-smooth. The rejection sampling method to sample from $\pi$ is the following:
  1. 1.
    Compute the minimizer $x^{*}$ of $V$, so that for any $Z$, $V(Z)\geq V(x^{*})+\frac{\beta}{2}\|Z-x^{*}\|^{2}$
  2. 2.
    Draw $Z\sim\mathcal{N}(x^{*},\frac{1}{\beta}I)$ and accept it with probability
    $\exp\left(-V(Z)+V(x^{*})+\frac{\beta}{2}\|Z-x^{*}\|^{2}\right).$
    Repeat this till acceptance.
The output of this method is distributed according to $\pi$ and the expected number of iterations is $(\frac{M}{\beta})^{d/2}$ [CGL+22, Theorem 7]. Applying this method to sample from $\tilde{\nu}_{y}$ gives a valid implementation of the RGO under smoothness of $f$ and appropriate step size with $\mathcal{O}(1)$ many iterations in expectation.
Appendix C Proof of Lemma 4
Proof.
Begin by recalling that if $X_{t}\sim\rho_{t}$ where $\,\mathrm{d}X_{t}=b_{t}(X_{t})\,\mathrm{d}t+\sqrt{2c}\,\mathrm{d}W_{t}$, then $\rho_{t}:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the Fokker-Planck PDE, given by:
$\partial_{t}\rho_{t}=-\nabla\cdot(\rho_{t}\,b_{t})+c\Delta\rho_{t}\,.$
Also note that since $\Delta\rho=\nabla\cdot(\rho\nabla\log\rho)$, the above can be written as:
$\partial_{t}\rho_{t}=-\nabla\cdot(\rho_{t}\,b_{t})+c\nabla\cdot(\rho_{t}\nabla\log\rho_{t})\,.$ (16)
We identify $\mu_{t}$ and $\nu_{t}$ with their densities with respect to Lebesgue measure, and further denote their relative density as $h_{t}=\frac{\mu_{t}}{\nu_{t}}$ We also assume enough regularity to take the differential under the integral sign and use $\int fg$ as a shorthand for $\int f(x)g(x)\,\mathrm{d}x$. Throughout the proof, we use integration by parts in various steps, denoted by (IBP). With all of this in mind, we have the following.
$\partial_{t}\,\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})$ $=\partial_{t}\int\nu_{t}\,\Phi(h_{t})$
$=\int(\partial_{t}\nu_{t})\Phi(h_{t})+\int\nu_{t}\,\big{(}\partial_{t}\,\Phi(h_{t})\big{)}$
$=\int(\partial_{t}\nu_{t})\Phi(h_{t})+\int\Phi^{\prime}(h_{t})\frac{\nu_{t}\,\partial_{t}\mu_{t}-\mu_{t}\,\partial_{t}\nu_{t}}{\nu_{t}^{2}}$
$=\underbrace{\int(\partial_{t}\nu_{t})\Phi(h_{t})}_{T_{1}}+\underbrace{\int\Phi^{\prime}(h_{t})(\partial_{t}\mu_{t})}_{T_{2}}-\underbrace{\int\Phi^{\prime}(h_{t})\frac{\mu_{t}}{\nu_{t}}(\partial_{t}\nu_{t})}_{T_{3}}$
We will now handle each of these terms separately. For the first term, we have:
$T_{1}$ $=\int(\partial_{t}\nu_{t})\Phi(h_{t})$
=(16)${{=}}\int\left(-\nabla\cdot(\nu_{t}\,b_{t})+c\nabla\cdot(\nu_{t}\nabla\log\nu_{t})\right)\Phi(h_{t})$ )
$=-\int\nabla\cdot(\nu_{t}\,b_{t})\Phi(h_{t})+c\int\nabla\cdot(\nu_{t}\nabla\log\nu_{t})\Phi(h_{t})$
=$\stackrel{{\scriptstyle\text{(IBP)}}}{{=}}\int\langle\nu_{t}b_{t},\nabla\Phi(h_{t})\rangle-c\int\langle\nu_{t}\nabla\log\nu_{t},\nabla\Phi(h_{t})\rangle$
$=\int\langle\nu_{t}b_{t},\Phi^{\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\rangle-c\int\langle\nu_{t}\nabla\log\nu_{t},\Phi^{\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\rangle$
We also have:
$T_{2}$ $=\int\Phi^{\prime}(h_{t})(\partial_{t}\mu_{t})$
=(16)${{=}}\int\Phi^{\prime}(h_{t})\left(-\nabla\cdot(\mu_{t}\,b_{t})+c\nabla\cdot(\mu_{t}\nabla\log\mu_{t})\right)$ )
=$\stackrel{{\scriptstyle\text{(IBP)}}}{{=}}\int\langle\mu_{t}\,b_{t},\nabla\Phi^{\prime}(h_{t})\rangle-c\int\langle\mu_{t}\nabla\log\mu_{t},\nabla(\Phi^{\prime}(h_{t}))\rangle$
$=\int\langle\mu_{t}\,b_{t},\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\rangle-c\int\langle\mu_{t}\nabla\log\mu_{t},\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\rangle$
We also have:
$T_{3}$ $=\int\Phi^{\prime}(h_{t})\frac{\mu_{t}}{\nu_{t}}(\partial_{t}\nu_{t})$
=(16)${{=}}\int\Phi^{\prime}(h_{t})\frac{\mu_{t}}{\nu_{t}}\left(-\nabla\cdot(\nu_{t}\,b_{t})+c\nabla\cdot(\nu_{t}\nabla\log\nu_{t})\right)$ )
=$\stackrel{{\scriptstyle\text{(IBP)}}}{{=}}\int\left\langle\nu_{t}\,b_{t},\nabla\left(\Phi^{\prime}(h_{t})\frac{\mu_{t}}{\nu_{t}}\right)\right\rangle-c\int\left\langle\nu_{t}\nabla\log\nu_{t},\nabla\left(\Phi^{\prime}(h_{t})\frac{\mu_{t}}{\nu_{t}}\right)\right\rangle$
$=\int\langle\mu_{t}\,b_{t},\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\rangle+\int\langle\nu_{t}b_{t},\Phi^{\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\rangle$
$\hskip 11.38092pt-c\int\left\langle\nabla\log\nu_{t},\mu_{t}\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\right\rangle-c\int\left\langle\nu_{t}\nabla\log\nu_{t},\Phi^{\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\right\rangle$
Therefore, combining the above, many terms cancel and we have the following:
$\partial_{t}\,\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})$ $=T_{1}+T_{2}-T_{3}$
$=c\int\left\langle\nabla\log\nu_{t},\mu_{t}\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\right\rangle-c\int\left\langle\mu_{t}\nabla\log\mu_{t},\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\right\rangle$
$=-c\int\left\langle\nabla\log\frac{\mu_{t}}{\nu_{t}},\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\right\rangle\mu_{t}$
$=-c\,\mathop{\mathbb{E}}_{\mu_{t}}\left[\left\langle\nabla\log\frac{\mu_{t}}{\nu_{t}},\Phi^{\prime\prime}(h_{t})\nabla\frac{\mu_{t}}{\nu_{t}}\right\rangle\right]=-c\,\mathsf{FI}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})\,,$
which proves the desired statement. ∎
Appendix D Proof of Lemma 5
Proof.
Let $\mu$ be an arbitrary probability distribution such that $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{\eta})<\infty$. As $\nu$ is $L$-smooth (i.e. $-LI\preceq\nabla^{2}f\preceq LI$) and $\eta\leq\frac{1}{L}$, $F(x)=x-\eta\nabla f(x)$ is a bijective map. Furthermore, it also holds that $\|F\|_{\mathsf{Lip}}\leq 1+\eta L$.
Therefore, from Lemma 1, we see that $\alpha_{\Phi\mathsf{SI}}(F_{\#}\nu^{\eta})\geq\frac{\alpha}{(1+\eta L)^{2}}$. Now for $t\geq 0$, let $\mu_{t}\coloneqq F_{\#}\mu*\mathcal{N}(0,tI)$ and $\nu_{t}\coloneqq F_{\#}\nu^{\eta}*\mathcal{N}(0,tI)$. For $t\geq 0$, further denote $\alpha_{\Phi\mathsf{SI}}(\nu_{t})=\alpha_{t}$ as shorthand. Also note that $\mathcal{N}(0,tI)$ is $1/t$-strongly log-concave and therefore from Lemma 3, $\alpha_{\Phi\mathsf{SI}}(\mathcal{N}(0,tI))\geq 1/t$ Hence, Lemma 2 implies that
$\alpha_{t}\geq\frac{\alpha}{(1+\eta L)^{2}+\alpha t}\,.$
The rate of change of $\Phi$-divergence between $\mu_{t}$ and $\nu_{t}$ is given by Lemma 4. Applying Lemma 4 for (10) with $b_{t}=0$ and $c=\frac{1}{2}$, which is what the (Gaussian convolution) evolution of $\mu_{t}$ and $\nu_{t}$ corresponds to, implies,
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})$ $=-\frac{1}{2}\mathsf{FI}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})$
$\leq-\alpha_{t}\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t}),$
where the inequality follows from $\nu_{t}$ satisfying a $\Phi$-Sobolev inequality. Applying Grönwall’s lemma and integrating the differential inequality from 00 to $2\eta$ yields,
$\frac{\mathsf{D}_{\Phi}(\mu_{2\eta}\mathbin{\|}\nu_{2\eta})}{\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})}\leq\exp\left(-\int_{0}^{2\eta}\alpha_{t}\,\mathrm{d}t\right)\leq\frac{(1+\eta L)^{2}}{(1+\eta L)^{2}+2\alpha\eta}\,.$
Finally, note that $\mathsf{D}_{\Phi}(\mu_{2\eta}\mathbin{\|}\nu_{2\eta})=\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}\mathbin{\|}\nu^{\eta}\bm{\mathrm{P}})$ and $\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})=\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{\eta})$, where the latter holds as $F$ is bijective. This, along with the fact that $\mu$ was arbitrary completes the proof. ∎
Appendix E Proof of Lemma 6
Proof.
Recall the ULA update in law (14). Under the assumptions of $\alpha$-strong log-concavity, $L$-smoothness, and $\eta\leq\frac{1}{L}$, it holds that $F(x)=x-\eta\nabla f(x)$ is a bijective map with $\|F\|_{\mathsf{Lip}}\leq 1-\alpha\eta$. Also note that $\mathcal{N}_{2\eta}$ is $\frac{1}{2\eta}$-SLC and therefore satisfies a $\Phi$-Sobolev inequality with the same constant (Lemma 3). Denote $\alpha_{\Phi\mathsf{SI}}(\nu^{\eta})=\beta$. Then Lemmas 1 and 2 imply that,
$\beta\geq\frac{1-(1-\alpha\eta)^{2}}{2\eta}=\frac{\alpha(2-\alpha\eta)}{2}\geq\frac{\alpha}{2}.$
Appendix F Deferred Proofs From Section 4
F.1 Proof of Lemma 8
Proof.
Let $\mu$ be an arbitrary distribution such that $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{X})<\infty$. Define $\mu_{t}\coloneqq\mu*N_{t}$ and $\nu_{t}\coloneqq\nu^{X}*\mathcal{N}_{t}$. Further denote $\alpha_{t}$ as shorthand for $\alpha_{\Phi\mathsf{SI}}(\nu_{t})$. By assumption, we know that $\nu^{X}$ satisfies a $\Phi$-Sobolev inequality with optimal constant $\alpha$. Also note that $\mathcal{N}_{t}$ is $\frac{1}{t}$-SLC and therefore satisfies a $\Phi$-Sobolev inequality with the same constant (Lemma 3). Therefore by Lemma 2 we have that
$\alpha_{t}\geq\frac{\alpha}{1+\alpha t}.$ (17)
Hence, applying Lemma 4 on the SDE $\,\mathrm{d}X_{t}=\,\mathrm{d}W_{t}$ gives us,
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{1}{2}\mathsf{FI}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})\,.$
We can then apply the $\Phi$-Sobolev inequality of $\nu_{t}$ and integrate the differential inequality from 00 to $\eta$ to yield
$\frac{\mathsf{D}_{\Phi}(\mu_{\eta}\mathbin{\|}\nu_{\eta})}{\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})}\leq\exp\left(-\int_{0}^{\eta}\alpha_{t}\,\mathrm{d}t\right)\leq\frac{1}{1+\alpha\eta}.$
Observe that $\mathsf{D}_{\Phi}(\mu_{\eta}\mathbin{\|}\nu_{\eta})=\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}^{+}\mathbin{\|}\nu^{X}\bm{\mathrm{P}}^{+})$ and that $\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})=\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{X})$. As $\mu$ was arbitrary, this gives a valid bound on the contraction coefficient and concludes the proof. ∎
F.2 Proof of Lemma 9
Proof.
The proof follows similarly to that of Lemma 8. Let $\mu$ be an arbitrary distribution such that $\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{Y})<\infty$. Define $\mu_{t}$ to be the marginal law at time $t$ when starting the SDE (15) from $\mu$, and let $\nu_{t}$ be the marginal law at time $t$ when starting the SDE (15) from $\nu^{Y}$. Further denote $\alpha_{t}$ as shorthand for $\alpha_{\Phi\mathsf{SI}}(\nu_{t})$. As by construction of the SDE (15), $\nu_{t}=\nu^{X}*\mathcal{N}_{\eta-t}$, the same argument used to derive (17) yields that
$\alpha_{t}\geq\frac{\alpha}{1+\alpha(\eta-t)}.$
Hence, applying Lemma 4 on the SDE (15) gives us,
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathsf{D}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})=-\frac{1}{2}\mathsf{FI}_{\Phi}(\mu_{t}\mathbin{\|}\nu_{t})\,.$
We can then apply the $\Phi$-Sobolev inequality of $\nu_{t}$ and integrate the differential inequality from 00 to $\eta$ to yield
$\frac{\mathsf{D}_{\Phi}(\mu_{\eta}\mathbin{\|}\nu_{\eta})}{\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})}\leq\exp\left(-\int_{0}^{\eta}\alpha_{t}\,\mathrm{d}t\right)\leq\frac{1}{1+\alpha\eta}.$
Observe that $\mathsf{D}_{\Phi}(\mu_{\eta}\mathbin{\|}\nu_{\eta})=\mathsf{D}_{\Phi}(\mu\bm{\mathrm{P}}^{-}\mathbin{\|}\nu^{Y}\bm{\mathrm{P}}^{-})$ and that $\mathsf{D}_{\Phi}(\mu_{0}\mathbin{\|}\nu_{0})=\mathsf{D}_{\Phi}(\mu\mathbin{\|}\nu^{Y})$. As $\mu$ was arbitrary, this gives a valid bound on the contraction coefficient and concludes the proof. ∎
Appendix G Ornstein-Uhlenbeck Process
Recall the Langevin dynamics (2) to sample from $\nu\propto e^{-f}$. When $\nu$ is a Gaussian, the dynamics are also called the Ornstein-Uhlenbeck process. For example, when $\nu=\mathcal{N}(0,\frac{1}{\alpha}I)$, the dynamics (2) become
$\,\mathrm{d}X_{t}=-\alpha X_{t}\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}W_{t}\,.$
In certain cases on the initial distribution $\rho_{0}$, the marginal distributions $\rho_{t}$ admit convenient Gaussian forms. We will leverage this to show that Theorems 1 and 2 are tight.
Throughout, let $\nu=\mathcal{N}(0,\frac{1}{\alpha}I)$ for some $\alpha>0$. Recall that this is $\alpha$-strongly log-concave, and hence also satisfies a $\Phi$-Sobolev inequality with the same constant (Lemma 3).
G.1 ULA evolution
The ULA update (3) for $\nu=\mathcal{N}(0,\frac{1}{\alpha}I)$ takes the following form:
$X_{k+1}=(1-\alpha\eta)X_{k}+\sqrt{2\eta}Z_{k}\,,$ (18)
and the solution to (18) is:
$X_{k}=(1-\alpha\eta)^{k}X_{0}+\sqrt{\frac{2(1-(1-\alpha\eta)^{2k})}{\alpha(2-\alpha\eta)}}Z\,,$ (19)
where $Z\sim\mathcal{N}(0,I)$ To show a tightness result, we take $\rho_{0}=\mathcal{N}(0,I)$ and $\Phi(x)=x\log x$ (i.e. Kl divergence) in the proposition below. In this case,
$\nu^{\eta}=\mathcal{N}\left(0,\frac{2}{\alpha(2-\alpha\eta)}I\right).$ (20)
Proposition 1.
Consider the ULA (3) for $\nu=\mathcal{N}(0,\frac{1}{\alpha}I)$ with $\rho_{0}=\mathcal{N}(0,I)$ and $\eta\leq\frac{1}{\alpha}$. Then,
$\mathsf{KL}(\rho_{k}\mathbin{\|}\nu^{\eta})=\frac{d}{2}\left[\frac{(1-\alpha\eta)^{2k}(2\alpha-\eta\alpha^{2}-2)}{2}-\log\left(1+\frac{(1-\alpha\eta)^{2k}(2\alpha-\eta\alpha^{2}-2)}{2}\right)\right].$ (21)
Therefore, $\mathsf{KL}(\rho_{k}\mathbin{\|}\nu^{\eta})=\tilde{\Theta}(d\alpha(1-\alpha\eta)^{2k})$.
Proof.
From (19), we can see that $X_{k}\sim\rho_{k}$ where
$\rho_{k}=\mathcal{N}\left(0,\frac{2+(1-\eta\alpha)^{2k}(2\alpha-\eta\alpha^{2}-2)}{\alpha(2-\eta\alpha)}I\right).$ (22)
Taking $k\to\infty$ in (22) we get that (20), i.e.
$\nu^{\eta}=\mathcal{N}\left(0,\frac{2}{\alpha(2-\alpha\eta)}I\right).$
Therefore, using the formula for KL divergence between two multivariate Gaussians, we get
$\mathsf{KL}(\rho_{k}\mathbin{\|}\nu^{\eta})=\frac{d}{2}\left[\frac{(1-\alpha\eta)^{2k}(2\alpha-\eta\alpha^{2}-2)}{2}-\log\left(1+\frac{(1-\alpha\eta)^{2k}(2\alpha-\eta\alpha^{2}-2)}{2}\right)\right].$
Remark 1.
Note that $\alpha$ refers to the assumption on $\nu$ in Proposition 1 whereas in Theorem 1 it refers to the $\Phi$-Sobolev constant of $\nu^{\eta}$. Therefore to meaningfully compare the two, $\alpha$ should be replaced by $\alpha(2-\alpha\eta)$ in (4) (following Lemma 6). Nonetheless, the tight exponential rate of convergence of Theorem 1 can be read off from (21).
G.2 Proximal sampler evolution
Similar to the ULA setting in Appendix G.1, we will take the target distribution to be $\nu^{X}=\mathcal{N}(0,\frac{1}{\alpha}I)$ and the starting distribution to be $\rho_{0}^{X}=\mathcal{N}(0,I)$. For this case, all of the iterates are Gaussian [CCSW22, Section 4.4]. Specifically, we denote $\rho_{k}^{X}=\mathcal{N}(0,c_{k}I)$ and $\rho_{k}^{Y}=\mathcal{N}(0,(c_{k}+\eta)I)$ where $c_{0}=1$. Then the proximal sampler (6) updates can be simplified to yield,
$c_{k+1}=\frac{c_{k}+\eta}{(1+\alpha\eta)^{2}}+\frac{\eta}{1+\alpha\eta}.$ (23)
We therefore have the following proposition.
Proposition 2.
Consider the proximal sampler (6) for $\nu^{X}=\mathcal{N}(0,\frac{1}{\alpha}I)$ with $\rho_{0}^{X}=\mathcal{N}(0,I)$ and $\eta>0$. Then,
$\mathsf{KL}(\rho_{k}^{X}\mathbin{\|}\nu^{X})=\frac{d}{2}\left[\frac{\alpha-1}{(1+\alpha\eta)^{2k}}-\log\left(1+\frac{\alpha-1}{(1+\alpha\eta)^{2k}}\right)\right].$
Therefore, $\mathsf{KL}(\rho_{k}^{X}\mathbin{\|}\nu^{X})=\tilde{\Theta}\left(d\alpha(1+\alpha\eta)^{-2k}\right)$.
Proof.
Solving (23), we get that
$c_{k}-1/\alpha=\frac{1-1/\alpha}{(1+\alpha\eta)^{2k}}$
Therefore, $\mathsf{KL}(\rho_{k}^{X}\mathbin{\|}\nu^{X})=\mathsf{KL}(\mathcal{N}(0,c_{k}I)\mathbin{\|}\mathcal{N}(0,\frac{1}{\alpha}I))$, which upon plugging in the formula for KL divergence between two multivariate Gaussians yields
$\frac{d}{2}\left[\frac{\alpha-1}{(1+\alpha\eta)^{2k}}-\log\left(1+\frac{\alpha-1}{(1+\alpha\eta)^{2k}}\right)\right].$
Conclusion
Therefore, from Corollary 1, Proposition 1, and Proposition 2, we can see that the rates of convergence along the continuous time Langevin dynamics, ULA, and the proximal sampler are $\left(d\alpha\exp(-2\alpha t)\right)$, $\left(d\alpha(1-\alpha\eta)^{2k}\right)$, and $\left(d\alpha(1+\alpha\eta)^{-2k}\right)$ respectively.